gaussian Elimination method

how to solve these equation using Gaussian Elimination Method with partial pivoting? 2x-7y+4z=9 x+9y-6z=1 -3x+8y+5z=6
Answer #1
2x-7y+4z=9 ------(1)
x+9y-6z=1------(2)
-3x+8y+5z=6 -----(3)
from (2) x=1-9y+6z
substituting x in (1) and (3)
2-18y+12z-7y+4z=9
-25y+16z=7-----(4)
-3+27y-18z+8y+5z=6
35y-13z=9-----(5)
solving (4),(5) to set y and z from this we get x
i hope this will help u
answered by: steve plz help reply